\providecommand{\href}[2]{#2}\begingroup\raggedright\begin{thebibliography}{10}

\bibitem{Policastro:2001yc}
G.~Policastro, D.~T. Son and A.~O. Starinets, \emph{{The Shear viscosity of
  strongly coupled N=4 supersymmetric Yang-Mills plasma}},
  \href{https://doi.org/10.1103/PhysRevLett.87.081601}{\emph{Phys. Rev. Lett.}
  {\bfseries 87} (2001) 081601}
  [\href{https://arxiv.org/abs/hep-th/0104066}{{\ttfamily hep-th/0104066}}].

\bibitem{Hartnoll:2016apf}
S.~A. Hartnoll, A.~Lucas and S.~Sachdev, \emph{{Holographic quantum matter}},
  \href{https://arxiv.org/abs/1612.07324}{{\ttfamily 1612.07324}}.

\bibitem{Grozdanov:2017ajz}
S.~Grozdanov, K.~Schalm and V.~Scopelliti, \emph{{Black hole scrambling from
  hydrodynamics}},
  \href{https://doi.org/10.1103/PhysRevLett.120.231601}{\emph{Phys. Rev. Lett.}
  {\bfseries 120} (2018) 231601}
  [\href{https://arxiv.org/abs/1710.00921}{{\ttfamily 1710.00921}}].

\bibitem{Blake:2018leo}
M.~Blake, R.~A. Davison, S.~Grozdanov and H.~Liu, \emph{{Many-body chaos and
  energy dynamics in holography}},
  \href{https://doi.org/10.1007/JHEP10(2018)035}{\emph{JHEP} {\bfseries 10}
  (2018) 035} [\href{https://arxiv.org/abs/1809.01169}{{\ttfamily
  1809.01169}}].

\bibitem{Maldacena:1997re}
J.~M. Maldacena, \emph{{The Large N limit of superconformal field theories and
  supergravity}}, \href{https://doi.org/10.1023/A:1026654312961}{\emph{Adv.
  Theor. Math. Phys.} {\bfseries 2} (1998) 231}
  [\href{https://arxiv.org/abs/hep-th/9711200}{{\ttfamily hep-th/9711200}}].

\bibitem{Gubser:1998bc}
S.~S. Gubser, I.~R. Klebanov and A.~M. Polyakov, \emph{{Gauge theory
  correlators from noncritical string theory}},
  \href{https://doi.org/10.1016/S0370-2693(98)00377-3}{\emph{Phys. Lett. B}
  {\bfseries 428} (1998) 105}
  [\href{https://arxiv.org/abs/hep-th/9802109}{{\ttfamily hep-th/9802109}}].

\bibitem{Witten:1998qj}
E.~Witten, \emph{{Anti-de Sitter space and holography}},
  \href{https://doi.org/10.4310/ATMP.1998.v2.n2.a2}{\emph{Adv. Theor. Math.
  Phys.} {\bfseries 2} (1998) 253}
  [\href{https://arxiv.org/abs/hep-th/9802150}{{\ttfamily hep-th/9802150}}].

\bibitem{Witten:1998zw}
E.~Witten, \emph{{Anti-de Sitter space, thermal phase transition, and
  confinement in gauge theories}},
  \href{https://doi.org/10.4310/ATMP.1998.v2.n3.a3}{\emph{Adv. Theor. Math.
  Phys.} {\bfseries 2} (1998) 505}
  [\href{https://arxiv.org/abs/hep-th/9803131}{{\ttfamily hep-th/9803131}}].

\bibitem{Son:2002sd}
D.~T. Son and A.~O. Starinets, \emph{{Minkowski space correlators in AdS / CFT
  correspondence: Recipe and applications}},
  \href{https://doi.org/10.1088/1126-6708/2002/09/042}{\emph{JHEP} {\bfseries
  09} (2002) 042} [\href{https://arxiv.org/abs/hep-th/0205051}{{\ttfamily
  hep-th/0205051}}].

\bibitem{Policastro:2002se}
G.~Policastro, D.~T. Son and A.~O. Starinets, \emph{{From AdS / CFT
  correspondence to hydrodynamics}},
  \href{https://doi.org/10.1088/1126-6708/2002/09/043}{\emph{JHEP} {\bfseries
  09} (2002) 043} [\href{https://arxiv.org/abs/hep-th/0205052}{{\ttfamily
  hep-th/0205052}}].

\bibitem{Nunez:2003eq}
A.~Nunez and A.~O. Starinets, \emph{{AdS / CFT correspondence, quasinormal
  modes, and thermal correlators in N=4 SYM}},
  \href{https://doi.org/10.1103/PhysRevD.67.124013}{\emph{Phys. Rev. D}
  {\bfseries 67} (2003) 124013}
  [\href{https://arxiv.org/abs/hep-th/0302026}{{\ttfamily hep-th/0302026}}].

\bibitem{Musiri:2003rs}
S.~Musiri and G.~Siopsis, \emph{{Asymptotic form of quasinormal modes of large
  AdS black holes}},
  \href{https://doi.org/10.1016/j.physletb.2003.10.015}{\emph{Phys. Lett. B}
  {\bfseries 576} (2003) 309}
  [\href{https://arxiv.org/abs/hep-th/0308196}{{\ttfamily hep-th/0308196}}].

\bibitem{BarraganAmado:2018zpa}
J.~Barrag\'an~Amado, B.~Carneiro Da~Cunha and E.~Pallante, \emph{{Scalar
  quasinormal modes of Kerr-AdS${_5}$}},
  \href{https://doi.org/10.1103/PhysRevD.99.105006}{\emph{Phys. Rev. D}
  {\bfseries 99} (2019) 105006}
  [\href{https://arxiv.org/abs/1812.08921}{{\ttfamily 1812.08921}}].

\bibitem{Amado:2021erf}
J.~B. Amado, B.~C. da~Cunha and E.~Pallante, \emph{{Quasinormal modes of scalar
  fields on small Reissner-Nordstr\"om-AdS5 black holes}},
  \href{https://doi.org/10.1103/PhysRevD.105.044028}{\emph{Phys. Rev. D}
  {\bfseries 105} (2022) 044028}
  [\href{https://arxiv.org/abs/2110.08349}{{\ttfamily 2110.08349}}].

\bibitem{Hortacsu:2011rr}
M.~Hortacsu, \emph{{Heun Functions and Some of Their Applications in Physics}},
   \href{https://arxiv.org/abs/1101.0471}{{\ttfamily 1101.0471}}.

\bibitem{fiziev2015heun}
P.~Fiziev, \emph{The heun functions as a modern powerful tool for research in
  different scientific domains}, {\emph{arXiv preprint arXiv:1512.04025} (2015)
  }.

\bibitem{ronveaux1995heun}
A.~Ronveaux and F.~M. Arscott, \emph{Heun's differential equations}. Oxford
  University Press, 1995.

\bibitem{Aminov:2020yma}
G.~Aminov, A.~Grassi and Y.~Hatsuda, \emph{{Black Hole Quasinormal Modes and
  Seiberg\textendash{}Witten Theory}},
  \href{https://doi.org/10.1007/s00023-021-01137-x}{\emph{Annales Henri
  Poincare} {\bfseries 23} (2022) 1951}
  [\href{https://arxiv.org/abs/2006.06111}{{\ttfamily 2006.06111}}].

\bibitem{Bianchi:2021xpr}
M.~Bianchi, D.~Consoli, A.~Grillo and J.~F. Morales, \emph{{QNMs of branes, BHs
  and fuzzballs from quantum SW geometries}},
  \href{https://doi.org/10.1016/j.physletb.2021.136837}{\emph{Phys. Lett. B}
  {\bfseries 824} (2022) 136837}
  [\href{https://arxiv.org/abs/2105.04245}{{\ttfamily 2105.04245}}].

\bibitem{Bonelli:2021uvf}
G.~Bonelli, C.~Iossa, D.~P. Lichtig and A.~Tanzini, \emph{{Exact solution of
  Kerr black hole perturbations via CFT2 and instanton counting: Greybody
  factor, quasinormal modes, and Love numbers}},
  \href{https://doi.org/10.1103/PhysRevD.105.044047}{\emph{Phys. Rev. D}
  {\bfseries 105} (2022) 044047}
  [\href{https://arxiv.org/abs/2105.04483}{{\ttfamily 2105.04483}}].

\bibitem{Bianchi:2021mft}
M.~Bianchi, D.~Consoli, A.~Grillo and J.~F. Morales, \emph{{More on the SW-QNM
  correspondence}}, \href{https://doi.org/10.1007/JHEP01(2022)024}{\emph{JHEP}
  {\bfseries 01} (2022) 024}
  [\href{https://arxiv.org/abs/2109.09804}{{\ttfamily 2109.09804}}].

\bibitem{Amado:2020zsr}
J.~B. Amado, B.~Carneiro~da Cunha and E.~Pallante, \emph{{Vector perturbations
  of Kerr-AdS$_{5}$ and the Painlev\'e VI transcendent}},
  \href{https://doi.org/10.1007/JHEP04(2020)155}{\emph{JHEP} {\bfseries 04}
  (2020) 155} [\href{https://arxiv.org/abs/2002.06108}{{\ttfamily
  2002.06108}}].

\bibitem{BarraganAmado:2021uyw}
J.~Barrag\'an~Amado, B.~Carneiro~da Cunha and E.~Pallante, \emph{{Remarks on
  holographic models of the Kerr-AdS$_{5}$ geometry}},
  \href{https://doi.org/10.1007/JHEP05(2021)251}{\emph{JHEP} {\bfseries 05}
  (2021) 251} [\href{https://arxiv.org/abs/2102.02657}{{\ttfamily
  2102.02657}}].

\bibitem{Novaes:2014lha}
F.~Novaes and B.~Carneiro~da Cunha, \emph{{Isomonodromy, Painlev\'e
  transcendents and scattering off of black holes}},
  \href{https://doi.org/10.1007/JHEP07(2014)132}{\emph{JHEP} {\bfseries 07}
  (2014) 132} [\href{https://arxiv.org/abs/1404.5188}{{\ttfamily 1404.5188}}].

\bibitem{Nekrasov:2009rc}
N.~A. Nekrasov and S.~L. Shatashvili, \emph{{Quantization of Integrable Systems
  and Four Dimensional Gauge Theories}},  in \emph{{16th International Congress
  on Mathematical Physics}}, pp.~265--289, 8, 2009,
  \href{https://arxiv.org/abs/0908.4052}{{\ttfamily 0908.4052}},
  \href{https://doi.org/10.1142/9789814304634_0015}{DOI}.

\bibitem{Alday:2009aq}
L.~F. Alday, D.~Gaiotto and Y.~Tachikawa, \emph{{Liouville Correlation
  Functions from Four-dimensional Gauge Theories}},
  \href{https://doi.org/10.1007/s11005-010-0369-5}{\emph{Lett. Math. Phys.}
  {\bfseries 91} (2010) 167} [\href{https://arxiv.org/abs/0906.3219}{{\ttfamily
  0906.3219}}].

\bibitem{Alday:2009fs}
L.~F. Alday, D.~Gaiotto, S.~Gukov, Y.~Tachikawa and H.~Verlinde, \emph{{Loop
  and surface operators in N=2 gauge theory and Liouville modular geometry}},
  \href{https://doi.org/10.1007/JHEP01(2010)113}{\emph{JHEP} {\bfseries 01}
  (2010) 113} [\href{https://arxiv.org/abs/0909.0945}{{\ttfamily 0909.0945}}].

\bibitem{Maruyoshi:2010iu}
K.~Maruyoshi and M.~Taki, \emph{{Deformed Prepotential, Quantum Integrable
  System and Liouville Field Theory}},
  \href{https://doi.org/10.1016/j.nuclphysb.2010.08.008}{\emph{Nucl. Phys. B}
  {\bfseries 841} (2010) 388}
  [\href{https://arxiv.org/abs/1006.4505}{{\ttfamily 1006.4505}}].

\bibitem{Belavin:1984vu}
A.~A. Belavin, A.~M. Polyakov and A.~B. Zamolodchikov, \emph{{Infinite
  Conformal Symmetry in Two-Dimensional Quantum Field Theory}},
  \href{https://doi.org/10.1016/0550-3213(84)90052-X}{\emph{Nucl. Phys. B}
  {\bfseries 241} (1984) 333}.

\bibitem{Piatek:2017fyn}
M.~Pi\k{a}tek and A.~R. Pietrykowski, \emph{{Solving Heun's equation using
  conformal blocks}},
  \href{https://doi.org/10.1016/j.nuclphysb.2018.11.021}{\emph{Nucl. Phys. B}
  {\bfseries 938} (2019) 543}
  [\href{https://arxiv.org/abs/1708.06135}{{\ttfamily 1708.06135}}].

\bibitem{Dolan:2000ut}
F.~A. Dolan and H.~Osborn, \emph{{Conformal four point functions and the
  operator product expansion}},
  \href{https://doi.org/10.1016/S0550-3213(01)00013-X}{\emph{Nucl. Phys. B}
  {\bfseries 599} (2001) 459}
  [\href{https://arxiv.org/abs/hep-th/0011040}{{\ttfamily hep-th/0011040}}].

\bibitem{Dolan:2003hv}
F.~A. Dolan and H.~Osborn, \emph{{Conformal partial waves and the operator
  product expansion}},
  \href{https://doi.org/10.1016/j.nuclphysb.2003.11.016}{\emph{Nucl. Phys. B}
  {\bfseries 678} (2004) 491}
  [\href{https://arxiv.org/abs/hep-th/0309180}{{\ttfamily hep-th/0309180}}].

\bibitem{Dolan:2011dv}
F.~A. Dolan and H.~Osborn, \emph{{Conformal Partial Waves: Further Mathematical
  Results}},  \href{https://arxiv.org/abs/1108.6194}{{\ttfamily 1108.6194}}.

\bibitem{Bonelli:2022ten}
G.~Bonelli, C.~Iossa, D.~P. Lichtig and A.~Tanzini, \emph{{Irregular Liouville
  correlators and connection formulae for Heun functions}},
  \href{https://arxiv.org/abs/2201.04491}{{\ttfamily 2201.04491}}.

\bibitem{Dodelson:2022yvn}
M.~Dodelson, A.~Grassi, C.~Iossa, D.~Panea~Lichtig and A.~Zhiboedov,
  \emph{{Holographic thermal correlators from supersymmetric instantons}},
  \href{https://arxiv.org/abs/2206.07720}{{\ttfamily 2206.07720}}.

\bibitem{Rodriguez-Gomez:2021pfh}
D.~Rodriguez-Gomez and J.~G. Russo, \emph{{Correlation functions in finite
  temperature CFT and black hole singularities}},
  \href{https://doi.org/10.1007/JHEP06(2021)048}{\emph{JHEP} {\bfseries 06}
  (2021) 048} [\href{https://arxiv.org/abs/2102.11891}{{\ttfamily
  2102.11891}}].

\bibitem{Rodriguez-Gomez:2021mkk}
D.~Rodriguez-Gomez and J.~G. Russo, \emph{{Thermal correlation functions in CFT
  and factorization}},
  \href{https://doi.org/10.1007/JHEP11(2021)049}{\emph{JHEP} {\bfseries 11}
  (2021) 049} [\href{https://arxiv.org/abs/2105.13909}{{\ttfamily
  2105.13909}}].

\bibitem{Georgiou:2022ekc}
G.~Georgiou and D.~Zoakos, \emph{{Holographic correlation functions at finite
  density and/or finite temperature}},
  \href{https://arxiv.org/abs/2209.14661}{{\ttfamily 2209.14661}}.

\bibitem{Matone:1995rx}
M.~Matone, \emph{{Instantons and recursion relations in N=2 SUSY gauge
  theory}}, \href{https://doi.org/10.1016/0370-2693(95)00920-G}{\emph{Phys.
  Lett. B} {\bfseries 357} (1995) 342}
  [\href{https://arxiv.org/abs/hep-th/9506102}{{\ttfamily hep-th/9506102}}].

\bibitem{MYERS1986304}
R.~Myers and M.~Perry, \emph{Black holes in higher dimensional space-times},
  \href{https://doi.org/https://doi.org/10.1016/0003-4916(86)90186-7}{\emph{Annals
  of Physics} {\bfseries 172} (1986) 304}.

\bibitem{Hawking:1998kw}
S.~W. Hawking, C.~J. Hunter and M.~Taylor, \emph{{Rotation and the AdS / CFT
  correspondence}},
  \href{https://doi.org/10.1103/PhysRevD.59.064005}{\emph{Phys. Rev. D}
  {\bfseries 59} (1999) 064005}
  [\href{https://arxiv.org/abs/hep-th/9811056}{{\ttfamily hep-th/9811056}}].

\bibitem{Tattersall:2018axd}
O.~J. Tattersall, \emph{{Kerr\textendash{}(anti\textendash{})de Sitter black
  holes: Perturbations and quasinormal modes in the slow rotation limit}},
  \href{https://doi.org/10.1103/PhysRevD.98.104013}{\emph{Phys. Rev. D}
  {\bfseries 98} (2018) 104013}
  [\href{https://arxiv.org/abs/1808.10758}{{\ttfamily 1808.10758}}].

\bibitem{Fitzpatrick:2012yx}
A.~L. Fitzpatrick, J.~Kaplan, D.~Poland and D.~Simmons-Duffin, \emph{{The
  Analytic Bootstrap and AdS Superhorizon Locality}},
  \href{https://doi.org/10.1007/JHEP12(2013)004}{\emph{JHEP} {\bfseries 12}
  (2013) 004} [\href{https://arxiv.org/abs/1212.3616}{{\ttfamily 1212.3616}}].

\bibitem{Komargodski:2012ek}
Z.~Komargodski and A.~Zhiboedov, \emph{{Convexity and Liberation at Large
  Spin}}, \href{https://doi.org/10.1007/JHEP11(2013)140}{\emph{JHEP} {\bfseries
  11} (2013) 140} [\href{https://arxiv.org/abs/1212.4103}{{\ttfamily
  1212.4103}}].

\bibitem{Kulaxizi:2019tkd}
M.~Kulaxizi, G.~S. Ng and A.~Parnachev, \emph{{Subleading Eikonal, AdS/CFT and
  Double Stress Tensors}},
  \href{https://doi.org/10.1007/JHEP10(2019)107}{\emph{JHEP} {\bfseries 10}
  (2019) 107} [\href{https://arxiv.org/abs/1907.00867}{{\ttfamily
  1907.00867}}].

\bibitem{Karlsson:2019dbd}
R.~Karlsson, M.~Kulaxizi, A.~Parnachev and P.~Tadi\'c, \emph{{Leading
  Multi-Stress Tensors and Conformal Bootstrap}},
  \href{https://doi.org/10.1007/JHEP01(2020)076}{\emph{JHEP} {\bfseries 01}
  (2020) 076} [\href{https://arxiv.org/abs/1909.05775}{{\ttfamily
  1909.05775}}].

\bibitem{Karlsson:2020ghx}
R.~Karlsson, M.~Kulaxizi, A.~Parnachev and P.~Tadi\'c, \emph{{Stress tensor
  sector of conformal correlators operators in the Regge limit}},
  \href{https://doi.org/10.1007/JHEP07(2020)019}{\emph{JHEP} {\bfseries 07}
  (2020) 019} [\href{https://arxiv.org/abs/2002.12254}{{\ttfamily
  2002.12254}}].

\bibitem{Li:2019zba}
Y.-Z. Li, \emph{{Heavy-light Bootstrap from Lorentzian Inversion Formula}},
  \href{https://doi.org/10.1007/JHEP07(2020)046}{\emph{JHEP} {\bfseries 07}
  (2020) 046} [\href{https://arxiv.org/abs/1910.06357}{{\ttfamily
  1910.06357}}].

\bibitem{Li:2020dqm}
Y.-Z. Li and H.-Y. Zhang, \emph{{More on heavy-light bootstrap up to
  double-stress-tensor}},
  \href{https://doi.org/10.1007/JHEP10(2020)055}{\emph{JHEP} {\bfseries 10}
  (2020) 055} [\href{https://arxiv.org/abs/2004.04758}{{\ttfamily
  2004.04758}}].

\end{thebibliography}\endgroup

\documentclass[12pt]{article}

\usepackage[normalem]{ulem}

\usepackage{xcolor}
\usepackage[english]{babel}
\usepackage[T1]{fontenc}
\usepackage[utf8]{inputenc}
\usepackage{authblk}
\usepackage{mathtools}
\usepackage{slashed}
\usepackage{amsmath,amssymb}
\usepackage{amsfonts}
\usepackage{enumitem}
\usepackage{graphicx,color,xcolor}
\usepackage{cite}
\usepackage{tabularx}
\usepackage{subcaption}
\usepackage{hyperref}
\hypersetup{
    colorlinks=true,
    linkcolor=blue,
    filecolor=magenta,      
    urlcolor=cyan,
}
\usepackage[capitalize]{cleveref}

\numberwithin{equation}{section} 

\definecolor{refcol}{rgb}{0.9,0.1,0.1}
\hypersetup{colorlinks=true,linkcolor=blue,citecolor=refcol,urlcolor=cyan,linktocpage}

\graphicspath{{Plots/}}

\begin{document}
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\begin{titlepage}
\thispagestyle{empty}

%\title{{\Huge\bf Exact thermal correlator of $CFT$ with chemical potential}}

\title{
{\Huge\bf Exact thermal correlators of holographic $CFT$s }
}

%\bigskip\bigskip\bigskip\bigskip\bigskip
\vfill

\author{
	{\bf Atanu Bhatta$^a$}\thanks{{\tt atanu.bhatta@wits.ac.za}},  
	{\bf Taniya Mandal$^a$}\thanks{{\tt taniya.mandal@wits.ac.za}}
	\smallskip\hfill\\      	
	\small{
		$^a${\it National Institute of Theoretical and Computational Sciences}\\
		{\it School of Physics and Mandelstam Institute for Theoretical Physics}\\
		{\it University of the Witwatersrand, Johannesburg, Wits 2050, South Africa.}
			}
		}

%\bigskip\bigskip\bigskip\bigskip
\vfill

\date{
\vspace{1cm}
\begin{quote}
\centerline{{\bf Abstract}}
{\small
We compute the exact retarded Green's functions in thermal $CFT$s with chemical potential and angular momenta using holography respectively. We consider the field equations satisfied by the quasi-normal modes in both charged and rotating black holes in $AdS$ spacetime and mapped them to the Heun equations by appropriate changes of variables. The AGT correspondence allows us to find the connection formulae among the solutions of the Heun equations near different singularities by using the crossing relations of the five-point correlators in the Liouville $CFT$. The connection formulae associated with the boundary conditions imposed on the bulk field equations yield the exact thermal correlators in the boundary $CFT$. 
}
\end{quote}
}

%\vfill
%\leftline{{\bf Report No: }} }

%\begin{abstract}
%\end{abstract}

\end{titlepage}
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\thispagestyle{empty}\maketitle\vfill \eject

\tableofcontents
%\newpage

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\section{Introduction}
The retarded Green's functions of a finite temperature conformal field theory ($CFT$) are useful to study quantum transports \cite{Policastro:2001yc,Hartnoll:2016apf}, chaotic behaviour \cite{Grozdanov:2017ajz,Blake:2018leo} etc. of the given system. The $AdS/CFT$ correspondence~\cite{Maldacena:1997re,Gubser:1998bc,Witten:1998qj} dictates that the dual gravity theory of a thermal $CFT$ residing at the boundary of an asymptotically $AdS$ space can be described by an $AdS$ black hole~\cite{Witten:1998zw}. In this setup, the retarded Green's functions of the boundary $CFT$ can be obtained by considering quasi-normal modes in the dual bulk with appropriate boundary conditions. Specifically, one imposes in-going boundary  conditions on the quasi-normal modes near the horizon and identifies the response and source terms of the solution at the asymptotic boundary. Then the ratio of the response function to the source yields the retarded Green's function~\cite{Son:2002sd,Policastro:2002se,Nunez:2003eq}.

The wave equation satisfied by the quasi-normal modes in the bulk can be recast into the Heun equation by making appropriate changes of variables and field redefinitions~\cite{Musiri:2003rs,BarraganAmado:2018zpa,Amado:2021erf,Hortacsu:2011rr}. The Heun equation is the linear second order differential equation with four regular singular points~\cite{fiziev2015heun,ronveaux1995heun}. One can solve the Heun equation around the singular points to get different solutions. These are basically the expansions of the Heun functions around the points under considerations. 

Quasi-normal modes of black holes can also be studied via quantum Seiberg-Witten curve for $\mathcal{N}=2$, SYM theory \cite{Aminov:2020yma,Bianchi:2021xpr,Bonelli:2021uvf,Bianchi:2021mft}. One also finds the Heun equations while studying the scattering of $Kerr-AdS$ black holes using monodromy\cite{Amado:2020zsr,BarraganAmado:2021uyw,Novaes:2014lha}. The Heun equation appears in the study of surface operators in $\mathcal{N}=2, SU(2)$ supersymmetric gauge theory with flavors $N_f \le 4$ in four dimensions \cite{Nekrasov:2009rc}. Moreover, the Alday-Gaiotto-Tachikawa (AGT) correspondence relates the BPS sector of the gauge theory in the so-called $\Omega$-background to the two-dimensional Liouville $CFT$ \cite{Alday:2009aq,Alday:2009fs}. The four-point conformal block of the primary fields in the Liouville $CFT$ corresponds to the Nekrasov partition function of the gauge theory with four flavors and inserting degenerate fields into the correlator amounts in inserting the surface operators in the gauge theory \cite{Maruyoshi:2010iu}. In particular, in the semi-classical limit the BPZ equation~\cite{Belavin:1984vu} satisfied by the five-point correlator with one degenerate insertion in the Liouville theory becomes the Heun equation, while the points of primary operator insertions are identified with the singular points of the Heun equation \cite{Piatek:2017fyn}. 

One can solve for the conformal blocks around the different points of primary operator insertions with appropriate boundary conditions~\cite{Dolan:2000ut,Dolan:2003hv,Dolan:2011dv}. These solutions are called the conformal blocks of different channels, namely $s, t$ and $u$ channels of the corresponding correlation function. The different channels are related to each other via crossing relations. Thus using crossing relations, the solution of the Heun equation near a given singular point admits an expansion in terms of the solutions around the other singularities. These are called connection formulas between the Heun functions around different singular points~\cite{Bonelli:2022ten}. 

Finally, from the connection formula of the Heun function associated with the in-going solutions of the bulk quasi-normal mode and that associated with the asymptotic behaviour of the quasi-normal mode one can read the source and the response function and hence obtain the retarded Green's function for the theory under consideration. In this way, a closed form expression of the exact Green's function has been obtained recently for the $AdS$-Schwarzschild background which describes a finite temperature $CFT$ at the boundary~\cite{Dodelson:2022yvn}.  Earlier, attempts to find thermal two-point correlators of holographic $CFT$s have been made using Geodesic approximation method, in the black brane background in the large conformal dimension limit  \cite{Rodriguez-Gomez:2021pfh,Rodriguez-Gomez:2021mkk,Georgiou:2022ekc}. Whereas the computation of the same using quasi-normal modes in the dual $AdS$ space and connecting with the Heun equation is exact \cite{Dodelson:2022yvn}.  

We want to generalize this idea for the charged black hole, namely the Reissner-Nordstr\"{o}m $AdS_5$ black hole, for which the boundary theory becomes a finite temperature $CFT_4$ with a chemical potential. The chemical potential is related to the charge of the black hole sitting in the bulk. We consider the charged scalar quasi-normal modes in the bulk, and obtain an expression for the exact Green's function following the prescription described above. We further extend the computation of exact thermal correlator for a $CFT_4$ that is dual to a rotational $AdS$ black hole.

Organisation of this paper is as follows: In section \ref{sec:rnbh}, we discuss the computation of the exact two-point thermal correlator of a $CFT$ with chemical potential. Retarded Greens function calculation for a rotational $CFT$ has been analyzed in section \ref{sec:rotbh}. Finally, we conclude in section \ref{sec:con}.

%
%\section{Charged $AdS_5$ black hole}\label{sec:rnbh}
%\section{Exact holographic thermal correlator of a $CFT$ dual to a charged $AdS_5$ black hole}\label{sec:rnbh}
\section{$CFT$ dual to a charged $AdS_5$ black hole}\label{sec:rnbh}
In this section we compute the exact thermal correlator of a $CFT$ that is dual to a five dimensional charged black hole in asymptotically Anti-de sitter space.
The five dimensional Einstein-Maxwell theory with a negative cosmological constant is described by the following action
\begin{eqnarray}
S =\frac{1}{16\pi G_5} \int d^5x \sqrt{-g} \left(R + 12- \frac{1}{g_s^2} F_{\mu\nu}F^{\mu\nu} \right)
\end{eqnarray}
%where we have set the $AdS$ radius $L=1$. 
Solution to the equations of motion derived from the above action is called the Reissner-Nordstr\"{o}m (RN) black hole. The five dimensional  Reissner-Nordstr\"{o}m black hole in $AdS$ space is given by  
\begin{eqnarray}
    ds^2 =-f(r)dt^2+\frac{dr^2}{f(r)}+r^2 d\Omega_3^2; \qquad A = \mu \left( 1- \frac{r_+^2}{r^2}\right) \,dt.
\end{eqnarray}
Whereas the blackening factor $f(r)$ and the chemical potential $\mu$ are  
\begin{eqnarray}
    f(r) = 1-\frac{M}{r^2}+\frac{Q^2}{r^4}+\frac{r^2}{L^2}, \quad \text{and} \quad \mu = \sqrt{\frac{3}{4}} \frac{g_s Q}{r_+^2},
\end{eqnarray}
with $r_+$ being the radius of the outer horizon. The black hole parameters $M$ and $Q$ are related to the ADM mass and charge of the black hole as
\begin{eqnarray}
 M_{ADM} = \frac{3 M \Omega_3}{16 \pi G_5}; \quad \text{and} \quad Q_e = \frac{\sqrt{3}Q \Omega_3}{4\pi G_5 g_s},
\end{eqnarray}
respectively. The radius of the outer horizon $r_+$ is related to the black hole parameters as
\begin{eqnarray}
 M = r_+^2 + \frac{Q^2}{r_+^2} + \frac{r_+^4}{L^2}.
\end{eqnarray}
In terms of the outer horizon $r_+$, the blackening factor $f(r)$ can be written as  \cite{Amado:2021erf}
\begin{eqnarray}
    f(r) =\frac{(r^2-r^2_+)(r^2-r^2_-)(r^2-r^2_0)}{L^2r^4},
\end{eqnarray}
where
\begin{eqnarray}\label{rmr0}
    r_-^2 &=& \frac{L^2}{2}\left(-1-\frac{r^2_+}{L^2} + \sqrt{\left(1+\frac{r^2_+}{L^2}\right)^2+\frac{4Q^2}{L^2 r^2_+}}\right),\nonumber\\
 r_0^2 &=& \frac{L^2}{2}\left(-1-\frac{r^2_+}{L^2} - \sqrt{\left(1+\frac{r^2_+}{L^2}\right)^2+\frac{4Q^2}{L^2 r^2_+}}\right).
 \end{eqnarray}
 Temperature at the outer horizon $r_+$ is given by 
 \begin{eqnarray}\label{temp}
  T = \frac{1}{\pi}\left(\frac{r_+}{L^2}+\frac{1}{2r_+}-\frac{Q^2}{2r_+^5}\right).
 \end{eqnarray}
 Here onward, for simplicity, we set the $AdS$ radius $L=1$. In this set up, we consider a charged scalar field excitation with mass $m$ and charge $e$ on the RN-$AdS$ black hole background. The mass of the bulk scalar field is related to the conformal dimension $\Delta$ of a scalar primary operator in the boundary $CFT$ as
 \begin{eqnarray}
 m^2 = \Delta(\Delta-4).
 \end{eqnarray}
 In the charged black hole background, the Klein-Gordon equation satisfied by the charged scalar field $\phi$ reads as
\begin{eqnarray}\label{cskg}
    \frac{1}{\sqrt{-g}}D_\mu\left(\sqrt{-g} g^{\mu\nu}D_\nu\right)\phi -m^2 \phi = 0,
\end{eqnarray}
where the covariant derivative $D_\mu =\nabla_\mu - i e A_\mu$. Decomposing the scalar field in the Fourier modes as follows
\begin{eqnarray}
    \phi(t,r,\Omega) = \int d\omega \sum_{\ell, m_1, m_2} e^{-i\omega t}Y_{\ell}^{m_1,m_2}(\Omega) \psi_{\omega \ell}(r),
\end{eqnarray}
we can write the wave equation \eqref{cskg} in the following form
\begin{eqnarray}
 \Bigg(   \frac{1}{r}\frac{d}{dr}\left(r^3 f(r)\frac{d}{dr}\right)+\frac{r^2}{f(r)}
    \left(\omega-\frac{\sqrt{3}eQ}{2r^2}\right)^2-m^2 r^2-\ell(\ell+2)\Bigg)\psi_{\omega \ell}(r)=0     .
\end{eqnarray}

We want to solve the wave equation with the following boundary conditions. At the outer horizon, $r = r_+$ we impose an in-going boundary condition as
\begin{eqnarray}
\label{eq:bc@horizon}
 \psi^{in}_{\omega \ell} (r) = (r-r_+)^{-i\frac{(\omega-\omega_c) r_+^3}{2(r_+^2-r_0^2)(r_+^2-r_-^2)}} + \cdots 
\end{eqnarray}
where,
\begin{eqnarray}
 \omega_c = \frac{\sqrt{3}e Q}{2r_+^2}.
\end{eqnarray}
The solution behaves at the boundary $(r \to \infty)$ as
\begin{eqnarray}
\label{eq:bc@bdy}
 \psi^{in}_{\omega \ell} (r) = \mathcal{A}(\omega, \ell) \left(r^{\Delta -4} + \cdots \right) +\mathcal{B}(\omega, \ell) \left(r^{-\Delta} + \cdots \right)  .
\end{eqnarray}
Then the retarded Green's function can be computed as
\begin{eqnarray}
 G_R(\omega, \ell) = \frac{\mathcal{B}(\omega, \ell)}{\mathcal{A}(\omega, \ell)}.
\end{eqnarray}
\subsection{Heun equation}
Now to compute the ratio of the respose $\mathcal{B}(\omega, \ell)$ to the source $\mathcal{A}(\omega, \ell)$, we follow the technique of~\cite{Dodelson:2022yvn}. 
We adopt the following variable change
\begin{eqnarray}
 z &=& \frac{r^2-r^2_-}{r^2-r_0^2}\\
 \psi_{\omega \ell}(r) &=& z^{-\frac{1}{2}} (z-z_0)^{-\frac{1}{2}}(1-z)^{1/2} \chi_{\omega \ell}(z)
\end{eqnarray}
and recast the wave equation to the Heun equation in normal form as
\begin{eqnarray}
\label{eq:heun}
\Bigg( \partial_z^2+\frac{\frac{1}{4}-a_1^2}{(z-1)^2}-\frac{\frac{1}{2}-a_0^2-a_1^2 a_z^2+a_{\infty}^2+u}{z(z-1)}+\frac{\frac{1}{4}-a_z^2}{(z-z_0)^2}+\frac{u}{z(z-z_0)}+\frac{\frac{1}{4}-a_0^2}{z^2}\Bigg) \chi_{\omega l}(z)=0.\nonumber\\
\end{eqnarray}
The Heun equation has four singularities at $z = 0, 1, z_0$ and $\infty$, where 
\begin{eqnarray}
\label{eq:z0}
z_0 = \frac{r^2_+-r^2_-}{r^2_+-r^2_0}.
\end{eqnarray}
The outer horizon $(r=r_+)$ and the boundary $(r \to \infty)$ are mapped to $z =z_0$ and $z=1$ respectively.

We identify the parameters of the Heun equation for the charged $AdS$ black hole to be
\begin{eqnarray}\label{heunchargepara}
a_0 &=& -\frac{i r_-(r_-^2 \omega-r_+^2 \omega_c)}{2(r_-^2-r_+^2)(1+2r_-^2+r_+^2)},\\
a_z &=& \frac{ir_+^3(\omega-\omega_c)}{2(r_+^2-r_-^2)(1+r_-^2+2r_+^2)},\\
a_1 &=& \frac{\Delta-2}{2},\\
a_\infty &=& \frac{\sqrt{(1+r_-^2+r_+^2)}\left((1+r_-^2+r_+^2)\omega+r_+^2\omega_c\right)}{2(1+2r_-^2+r_+^2)(1+r_-^2+2r_+^2)},
\end{eqnarray}
and 
\begin{eqnarray}
\label{eq:u}
u &=& -\frac{l(l+2)+2(1+r_-^2+2r_+^2)+r_+^2\Delta(\Delta-4)}{4(1+2r_-^2+r_+^2)}\nonumber\\
&&+\frac{r_+^4(\omega-\omega_c)\left(r_-^2(\omega_c-3\omega)+r_+^2(\omega+\omega_c)\right)}{4(r_-^2-r_+^2)^2(1+2r_-^2+r_+^2)(1+r_-^2+2r_+^2)}.
\end{eqnarray}
 In the uncharged limit $Q\rightarrow 0\implies r_-\rightarrow 0, \omega_c\rightarrow 0$, the parameters \eqref{heunchargepara}-\eqref{eq:u} boil down to \cite{Dodelson:2022yvn}
\begin{eqnarray}\label{unchargedheunchargepara1}
	\tilde{a}_0 &=& 0,\\\label{unchargedheunchargepara2}
	\tilde{a}_z &=& \frac{ir_+\omega}{2(1+2r_+^2)},\\\label{unchargedheunchargepara3}
	\tilde{a}_1 &=& \frac{\Delta-2}{2},\\\label{unchargedheunchargepara4}
	\tilde{a}_\infty &=& \frac{\sqrt{(1+r_+^2)}\omega}{2(1+2r_+^2)},\\\label{unchargedheunchargepara5}
	\tilde{u} &=& -\frac{l(l+2)+2(1+2r_+^2)+r_+^2\Delta(\Delta-4)}{4(1+r_+^2)}\nonumber\\
	&&+\frac{r_+^2\omega ^2}{4(1+r_+^2)(1+2r_+^2)}.
\end{eqnarray}

The boundary conditions~\eqref{eq:bc@horizon},~\eqref{eq:bc@bdy} for the in-going solution take the following form at the horizon as
\begin{eqnarray}
\label{eq:zbc@horizon}
\chi_{\omega \ell}^{(z_0)} (z) = (z_0 - z)^{\frac{1}{2}-a_z} + \cdots 
\end{eqnarray}
and at the boundary as
\begin{eqnarray}
\label{eq:zbc@bdy}
\chi_{\omega \ell}^{(1)} (z)\propto \mathcal{A}(\omega, \ell) \chi^{(1), -}_{\omega \ell} (z) +\mathcal{B}(\omega, \ell) \chi^{(1), +}_{\omega \ell} (z)
\end{eqnarray}
where
\begin{eqnarray}
 \chi^{(1), \pm}_{\omega \ell} (z) = (1-z)^{1/2 \pm a_1} + \cdots 
\end{eqnarray}
\subsection{Relation to Liouville $CFT$}
The Heun equation appears in the study of conformal blocks in the Liouville $CFT$. The five-point correlation function with one degenerate operator insertion in Liouville $CFT$ satisfies the following BPZ equation 
\begin{eqnarray}
&&\left(b^{-2} \partial_z^2 + \frac{\Delta_1}{(z-1)^2} - \frac{\Delta_1 + z_0 \partial_{z_0} + \Delta_{z_0} + z\partial_z + \Delta_{2,1}+ \Delta_0 -\Delta_\infty }{z(z-1)} +\frac{\Delta_{z_0}}{(z-z_0)^2} \right. \cr 
 && \qquad \qquad \qquad + \left. \frac{z_0}{z(z-z_0)}\partial_{z_0}-\frac{1}{z}\partial_z + \frac{\Delta_0}{z^2}\right) \langle \Delta_{\infty} |V_1(1) V_{z_0}(z_0) \Phi(z) |\Delta_0 \rangle =0,
 \label{eq:bpz}
\end{eqnarray}
where $\Delta_1$ and $\Delta_{z_0}$ are the scaling dimensions of the operators $V_1$ and $V_{z_0}$ respectively and the scaling dimension of the degenerate operator $\Phi$ is given by $\Delta_{2,1} = -\frac{1}{2} - \frac{3b^2}{4}$. Since the conformal correlators admits an expansion in terms of the conformal blocks, the five-point conformal blocks also satisfy the BPZ equation. The ODE~\eqref{eq:bpz} has four regular singularities at $z=0, 1, z_0$ and $\infty$. One can solve the ODE around different singular points and obtain different five-point blocks using appropriate boundary conditions. These answers for the five-point blocks are related to each other by crossing relations.

Now we take the semi-classical limit,
\begin{eqnarray}
b \to 0, \alpha_i \to 0, b \alpha_i = a_i (\text{finite}),
\end{eqnarray}
where the momenta $\alpha_i$'s of the primary operators are related to the scaling dimensions as
\begin{eqnarray}
\Delta_i = \alpha_i (Q- \alpha_i); \qquad Q = \left(b + \frac{1}{b} \right).
\end{eqnarray}
In the semi-classical limit, the BPZ equation~\eqref{eq:bpz} reduces to the Heun equation in normal form~\eqref{eq:heun}. The operator insertion points $z =0, 1, z_0$ and $\infty$ becomes the four singular points of corresponding Heun equation and the crossing relations among the conformal blocks descend to the connection formula between the solutions of the Heun equation.

The connection formula between the solution~\eqref{eq:zbc@horizon} near the horizon $(z = z_0)$ and the solution~\eqref{eq:zbc@bdy} near the boundary $z =1$ reads~\cite{Bonelli:2022ten,Dodelson:2022yvn}
\begin{eqnarray}
\label{eq:connection}
\chi^{(z_0)}_{\omega \ell} (z) &=& \sum_{\theta'=\pm} \left(\sum_{\theta=\pm} \mathcal{M}_{-\theta}(a_{z}, a; a_0)\mathcal{M}_{(-\theta)\theta'}(a, a_1; a_{\infty}) \right. \cr  
&& \qquad \qquad \left. z_0^{\theta a}e^{-\frac{\theta}{2}\partial_a F} \right) z_0^{\frac{1}{2}-a_0-a_{z}}(1-z_0)^{a_z-a_1} e^{-\frac{1}{2}(\partial_{a_z}+\theta' \partial_{a_1})F} \chi_{\omega \ell}^{(1), \theta'}(z)\nonumber, \\
\end{eqnarray}
with
\begin{eqnarray}
\mathcal{M}_{\theta\theta'} (a_0, a_1; a_2) &=& \frac{\Gamma(-2\theta' a_1)}{\Gamma(\frac{1}{2} + \theta a_0- \theta' a_1 + a_2)} \frac{\Gamma(1+2\theta a_0)}{\Gamma(\frac{1}{2} + \theta a_0- \theta' a_1 - a_2)}.
\end{eqnarray}
Where $\theta=\pm$ are the two fusion channels allowed by the degenerate fusion rules, $a$ is defined by the Matone relation \cite{Matone:1995rx}
\begin{eqnarray}
\label{eq:matone}
u = -a^2 + a_z^2 -\frac{1}{4} + a_0^2 + z_0 \partial_{z_0} F,
\end{eqnarray}
and $F(z_0)$ is the classical conformal block, related to the conformal block without degenerate insertion as
\begin{eqnarray}
\mathcal{F}(\Delta_{0}, \Delta_1, \Delta_{z_0}, \Delta_{\infty}; \Delta; z_0) = z_0^{\Delta-\Delta_{z_0}-\Delta_0} e^{b^{-2}(F(z_0) + \mathcal{O}(b^2)} .
\end{eqnarray}
\subsection{Retarded Green's function}
We can read the ratio of the response to the source from the connection formula~\eqref{eq:connection} and write the exact retarded Green's function as 
\begin{eqnarray}
 G_R(\omega, \ell) = (1 + r_+^2)^{2a_1} e^{-\partial_{a_1}F} \frac{\sum_{\theta=\pm} \mathcal{M}_{-\theta}(a_{z}, a; a_0)\mathcal{M}_{(-\theta)+}(a, a_1; a_{\infty})  z_0^{\theta a}e^{-\frac{\theta}{2}\partial_a F} }{\sum_{\theta=\pm} \mathcal{M}_{-\theta}(a_{z}, a; a_0)\mathcal{M}_{(-\theta)-}(a, a_1; a_{\infty})  z_0^{\theta a}e^{-\frac{\theta}{2}\partial_a F}}
\end{eqnarray}
where the parameters $z_0, a_0, a_1, a_z, a_\infty$ and $u$ are written in terms of $\omega, \ell$ and the RN-$AdS$ black hole parameters $r_0$ and $r_-$ as in~\eqref{eq:z0}-\eqref{eq:u}. One can compute $a$ using the Matone relation~\eqref{eq:matone}.

\subsection{Large black hole limit}
In this section, we are considering a case of large RN-$AdS$ black hole i.e. the outer horizon radius is very much greater than the $AdS$ length, $r_+\gg L$. From \eqref{rmr0}, we see that in this limit $r_-\rightarrow 0$ whereas $r_0^2\rightarrow -(1+r_+^2)$. Then for the large black holes, the parameters \eqref{heunchargepara}-\eqref{eq:u} of the Heun equation become 
\begin{eqnarray}\label{heunlarge}
 a_0 &=& 0\\
 a_z &=& \frac{i r_+(\omega-\omega_c)}{2(1+2r_+^2)}\\
 a_1 &=& \frac{\Delta-2}{2}\\
 a_{\infty} &=& \frac{\omega(1+r_+^2)+r_+^2\omega_c}{2\sqrt{(1+r_+^2)}(1+2r_+^2)}\\\label{eq:largeu}
 u &=& -\frac{l(l+2)+2(1+2r_+^2)+r_+^2\Delta(\Delta-4)}{4(1+r_+^2)}+\frac{r_+^2(\omega^2-\omega_c^2)}{4(1+r_+^2)(1+2r_+^2)}
\end{eqnarray}

In this case, we can express the above parameters \eqref{heunlarge}-\eqref{eq:largeu} in terms of uncharged parameters \eqref{unchargedheunchargepara1}- \eqref{unchargedheunchargepara5} as follows

\begin{eqnarray}\label{correctedheunlarge}
	a_0 &=& \tilde{a}_0\\
	a_z &=& \tilde{a}_z-\frac{i r_+\omega_c}{2(1+2r_+^2)}\\
	a_1 &=& \tilde{a}_1\\
	a_{\infty} &=&\tilde{a}_\infty+\frac{r_+^2\omega_c}{2\sqrt{(1+r_+^2)}(1+2r_+^2)}\\
	u &=& \tilde{u}-\frac{r_+^2\omega_c^2}{4(1+r_+^2)(1+2r_+^2)}
\end{eqnarray}
where $\omega_c=\frac{\sqrt{3}e Q}{2r_+^2}$.

%As, in our case, $L=1$, 
For a large black hole with small chemical potential $\mu\ll1$, $\omega_c$ becomes negligible and the temperature \eqref{temp} becomes
\begin{eqnarray}\label{largetemp}
  T = \frac{r_+}{\pi}.
 \end{eqnarray}
As $r_+\gg 1$, we find a high temperature black hole in this limit. In the limit $r_+\gg 1, \mu\ll 1$, the parameters of the Heun equation are

\begin{eqnarray}\label{heunlargesmallmu}
 a_0 &=& 0\nonumber\\
 a_z &=& \frac{i r_+\omega}{2(1+2r_+^2)}\nonumber\\
 a_1 &=& \frac{\Delta-2}{2}\nonumber\\
 a_{\infty} &=& \frac{\omega\sqrt{(1+r_+^2)}}{2(1+2r_+^2)}\nonumber\\
u &=& -\frac{l(l+2)+2(1+2r_+^2)+r_+^2\Delta(\Delta-4)}{4(1+r_+^2)}+\frac{r_+^2\omega^2}{4(1+r_+^2)(1+2r_+^2)}
\end{eqnarray}
Interestingly, from \eqref{heunlargesmallmu}, one notices that the parameters of the Heun equations of the charged, large black holes in the small chemical potential, high temperature limit are same as of uncharged black holes \cite{Dodelson:2022yvn}.

%\section{Rotational $AdS$ black hole} \label{sec:rotbh}
%\section{Exact holographic thermal correlator of a $CFT$ dual to a rotational $AdS_5$ black hole}\label{sec:rotbh}
\section{ $CFT$ dual to a rotational $AdS_5$ black hole}\label{sec:rotbh}
In this section we compute the exact, thermal two-point correlator of a $CFT$ dual to a rotating black hole in $AdS_5$ in the slow rotation approximation. In particular, we consider Myers-Perry black hole solution~\cite{MYERS1986304} in $AdS$ with a single rotation parameter $\alpha$ \cite{Hawking:1998kw}. We take into account the slow rotation i.e. $\alpha \ll 1$ and the solution is linear in $\alpha$~\cite{Tattersall:2018axd}. In this limit the metric is
\begin{eqnarray}
 ds^2& =& -\left(1-\frac{2M}{r^2}+r^2\right)dt^2+\frac{dr^2}{1-\frac{2M}{r^2}+r^2}+r^2 d\theta^2+r^2 \sin^2{\theta} d\phi^2+r^2 \cos^2{\theta} d\psi^2\nonumber\\&&-2\alpha\left(\frac{2M}{r^2}-r^2\right)\sin^2{\theta}dt d\phi.
\end{eqnarray}
Temperature  and the angular velocity of the slow rotating, one-parameter rotational black hole are given by
\begin{eqnarray}
 T=\frac{1+2r_+^2}{2\pi r_+},\quad\Omega =\frac{\alpha}{r_+^2},
\end{eqnarray}
where $r_+$ is the horizon and $AdS$ length has been set to $1$. 

We aim to solve the Klein-Gordon equation of a scalar field $\Phi$ of mass $\mu$ in this background. The equation is
\begin{eqnarray}
 \frac{1}{\sqrt{-g}}\partial_\mu (\sqrt{-g} g^{\mu\nu}\partial_\nu)\Phi-\mu^2\Phi=0.
\end{eqnarray}
Expanding the field in Fourier basis, in the linear order of $\alpha$, we can write the scalar field as \cite{BarraganAmado:2018zpa}
\begin{equation}
    \Phi = \int d\omega \sum_{l,m_1,m_2} e^{-i \omega t}e^{i m_1 \phi}e^{i m_2 \psi}\mathcal{P}(\theta)\Psi_{\omega }(r).
\end{equation}
Then the scalar wave equation takes the following form
\begin{eqnarray}
&&\frac{\mathcal{P}(\theta)}{r^3}\partial_r\left(r^3 f(r)\partial_r\right)\Psi_{\omega l}(r)+\Bigg(\frac{\omega^2}{f(r)}+\frac{2\alpha m_1 \omega(1-\frac{2M}{r^4})}{f(r)}\Bigg)\mathcal{P}(\theta)\Psi_{\omega l}(r)-\frac{m_1^2}{r^2\sin^2\theta}\mathcal{P}(\theta)\Psi_{\omega l}(r)\nonumber\\
&&+\Bigg(\frac{1}{\sin( 2\theta)}\partial_\theta(\sin (2\theta)\partial_\theta)\mathcal{P}(\theta)-\frac{m_2^2}{\cos^2\theta}\mathcal{P}(\theta)\Bigg)\frac{\Psi_{\omega l}(r)}{r^2}-\mu^2 \Psi_{\omega l}(r)\mathcal{P}(\theta)=0
\end{eqnarray}
where $f(r)=1-\frac{2M}{r^2}+r^2$. Using separation of variable method, we find a radial and an angular equation as follows
\begin{eqnarray}\label{radial}
\frac{1}{r^3}\partial_r\left(r^3 f(r)\partial_r\right)\Psi_{\omega l}(r)+\Bigg(\frac{\omega^2}{f(r)}+\frac{2\alpha m_1 \omega(1-\frac{2M}{r^4})}{f(r)}\Bigg)\Psi_{\omega }(r) -\lambda \frac{\Psi_{\omega }(r)}{r^2}-\mu^2 \Psi_{\omega }(r)=0\nonumber\\
\end{eqnarray}
\begin{eqnarray}\label{angular}
m_1^2\mathcal{P}(\theta)-\sin^2\theta\Bigg(\frac{1}{\sin( 2\theta)}\partial_\theta(\sin (2\theta)\partial_\theta)\mathcal{P}(\theta)-\frac{m_2^2}{\cos^2\theta}\mathcal{P}(\theta)\Bigg)-\lambda \sin^2\theta \mathcal{P}(\theta)=0
\end{eqnarray}
Here $\lambda$ is the separation constant and the angular equation \eqref{angular} can be solved using Associated Legendre Polynomial. For example, for a special case when $m_1=m_2=m$, \eqref{angular} will be solved by the Associated Legendre Polynomial $P_l^m(\cos{(2\theta)}$ with $\lambda =4 l(l+1)$.
\subsection{Heun equation}
Under the following changes of variables
\begin{eqnarray}
z &=& \frac{r}{r^2+r_+^2+1},\\
\Psi_{\omega} (r) &=& \left(r^3 f(r)\frac{dz}{dr}\right)^{-1/2}\mathcal{R}_{\omega}(z),
\end{eqnarray}
the radial equation \eqref{radial} can be recast as a Heun equation
\begin{eqnarray}
\label{eq:rotatingheun}
\Bigg( \partial_z^2+\frac{\frac{1}{4}-a_1^2}{(z-1)^2}-\frac{\frac{1}{2}-a_0^2-a_1^2 a_z^2+a_{\infty}^2+u}{z(z-1)}+\frac{\frac{1}{4}-a_z^2}{(z-z_0)^2}+\frac{u}{z(z-z_0)}+\frac{\frac{1}{4}-a_0^2}{z^2}\Bigg) \mathcal{R}_{\omega }(z)=0.\nonumber\\
\end{eqnarray}
Where the parameters are
\begin{eqnarray}
\label{eq:rota0}
a_0 &=& 0,\\
a_1 &=& \frac{\Delta-2}{2} ,\\
a_z &=& i\frac{\omega}{2}\frac{r_+\sqrt{1-\frac{2\alpha m_1}{\omega r_+^2}}}{1+2r_+^2},\\
a_{\infty} &=& \frac{\omega}{2}\frac{\sqrt{1+r_+^2+\frac{2\alpha m_1}{\omega}}}{1+2r_+^2},
\end{eqnarray}
\begin{eqnarray}
u = -\frac{\lambda + 2(2r_+^2+1)+r_+^2\Delta(\Delta-4)}{4(1+r_+^2)}+\frac{\alpha m_1 \omega}{2(1+r_+^2)}+\frac{r_+^2}{1+r_+^2}\frac{\omega^2}{4(2r_+^2+1)},
\end{eqnarray}
and the horizon $r_+$ is mapped to 
\begin{eqnarray}
\label{eq:rotz0}
z_0 &=& \frac{r_+^2}{2r_+^2+1}.
\end{eqnarray}
In the large black hole limit $r_+\gg 1$, if the angular velocity is small enough such that $\Omega\ll1$, the above parameters are same as those of Schwarzschild black hole when $\lambda= l(l+2)$. 
The Heun equation~\eqref{eq:rotatingheun} has four regular singularites at $z=0, 1, z_0$ and $\infty$.
\subsection{Retarded Green's function}
Following the same procedure as we did in the last section for the charged black hole, we compute the exact retarded Green's function for the rotating case as
\begin{eqnarray}
 G_R(\omega, \ell) = (1 + r_+^2)^{2a_1} e^{-\partial_{a_1}F} \frac{\sum_{\theta=\pm} \mathcal{M}_{-\theta}(a_{z}, a; a_0)\mathcal{M}_{(-\theta)+}(a, a_1; a_{\infty})  z_0^{\theta a}e^{-\frac{\theta}{2}\partial_a F} }{\sum_{\theta=\pm} \mathcal{M}_{-\theta}(a_{z}, a; a_0)\mathcal{M}_{(-\theta)-}(a, a_1; a_{\infty})  z_0^{\theta a}e^{-\frac{\theta}{2}\partial_a F}}
\end{eqnarray}
where the parameters $z_0, a_0, a_1, a_z, a_\infty$ and $u$ are written in terms of $\omega, m_1, \lambda$ and the black hole parameters $r_+$ and $\alpha$ as in~\eqref{eq:rota0}-\eqref{eq:rotz0}.
\section{Conclusion}\label{sec:con}

We have computed exact retarded Green's functions in the thermal $CFT$ with a chemical potential and that in the thermal $CFT$ with angular momenta holographically by considering Reissner-Nordstr\"{o}m and Myers-Perry black hole solutions in $AdS$ as the bulk theories respectively. We mapped the bulk field equations to the Heun equations by making appropriate changes of variables and a field redefinition. The Heun equation is mapped to the semi-classical BPZ equation satisfied by the five point correlators with degenerate insertion in the Liouville $CFT$. Using the crossing relations of the $CFT$ correlators we find the connection formulae for the solutions of the Heun equation of our interest near the regular singularities. From this connection formulae we read out the source and the response function for the bulk field and using the holographic prescription~\cite{Son:2002sd} we obtain the retarded Green's function for the scalar primary operators in the 4-dimensional thermal $CFT$.

The obvious next step would be to compute higher-point correlation functions in thermal $CFT$s. It would be interesting if this technique can be used to compute the out-of-time-ordered-correlators ($OTOC$) in the thermal $CFT$s. So far the exact retarded Green's functions are obtained for the 4-dimensional thermal $CFT$s. It would be very interesting if we can generalise the formalism to the d-dimensional case. 

We would like to compute current-current correlators, particularly $\langle J_\mu J_\nu \rangle$ and $\langle T_{\mu\nu} T_{\rho\sigma} \rangle$ in the thermal $CFT$s using this technique in future. These correlators are very useful to study transport phenomena and hydrodynamics.

We would like to interpret our results for two-point retarded Green's functions computed in the RN-$AdS$ and the rotating $AdS$ black hole backgrounds as heavy-light four point functions of zero temperature $CFT$ and solve for the heavy-light light-cone conformal bootstrap~\cite{Bonelli:2022ten,Fitzpatrick:2012yx,Komargodski:2012ek,Kulaxizi:2019tkd,Karlsson:2019dbd,Karlsson:2020ghx,Li:2019zba,Li:2020dqm}. It will also be very interesting to explore more in this direction to understand the underlying relations of 2d Liouville $CFT$ with the higher dimensional thermal $CFT$ precisely.

\section*{Acknowledgments} The work of AB is supported by the National Institute for Theoretical and Computational Sciences, NRF Grant Number 65212. The work of TM is supported by a Simons Foundation Grant Award ID 509116 and by the South African Research Chairs initiative of the Department of Science and Technology and the National Research Foundation.
\appendix 

\bibliographystyle{jhep}
\bibliography{bib_thermal}

\end{document}